How many liters of pure water must be mixed with 1 liter of a 5% vinegar solution to give a 2% solution?
Found 2 solutions by macston, Alan3354:
Answer by macston(5194)

(0.05)(1 liter)=0.05 liter of vinegar One liter of %5 solution contains 0.05 liter vinegar.
0.05 l/0.02=2.5 liters We have enough vinegar to make a total of 2.5 liters of 2% solution. We already have 1 liter so 2.5 liter-1 liter=1.5 liter We need to add 1.5 liters of water to dilute the concentration to 2%.

Answer by Alan3354(69443)

How many liters of pure water must be mixed with 1 liter of a 5% vinegar solution to give a 2% solution?
----------
Add W liters of water.
1*5 + 0*W = (1 + W)*2
5 = 2W + 2
W = 1.5 liters